OpenStudy (anonymous):

|7x + 7| = 56 A. x = 7 or x = -9 B. x = -7 or x = 9 C. x = 5 or x = -8 D.x = -5 or x = 8

sam (.sam.):

|7x + 7| = 56 \[(7x+7)^2=56^2 \\ \\ \text{Square both sides,} \\ \\ 49x^2+98x+49=3136 \\ \\ 49x^2+98x-3087=0 \\ \\ \text{Divide by 49,} \\ \\ x^2+2x-63=0 \\ \\ \text{factor,} \\ \\ (x-7)(x+9)=0 \\ \\ \text{Solve}\]

OpenStudy (anonymous):

idk how to do this math im home schooled and in 5th i cant do this math

sam (.sam.):

Where you don't understand?